# FLARE: Fast Low-rank Attention Routing Engine

### v1

@@ 5 Conclusion
FLARE is a token mixing layer that bypasses the quadratic cost of self-attention by leveraging low-rankness. Mechanically, FLARE routes attention through a fixed-size latent sequence via cross-attention projection and unprojection. FLARE achieves SOTA accuracy on a set of diverse PDE benchmarks, and comfortably scales to million-scale meshes and comfortably scales to PDE problems with million-scale geometries.
As transformers are the backbone of modern deep learning, we postulate that an efficient attention mechanism has several applications. We also identify potential areas for improvement: FLARE’s reliance on deep residual MLPs can introduce sequential bottlenecks and increase latency, suggesting that further speedups are possible by addressing this issue. Additional enhancements include (1) incrementally increasing the number of latent tokens during training; and (2) conditioning latent tokens on time for diffusion modeling.
@@ C.2 Benchmark datasets
The DrivAerML dataset (Ashton et al., 2024) has high-fidelity automotive aerodynamic simulations, featuring 500 parametrically morphed DrivAer vehicles. CFD simulations are performed on 160 million volumetric mesh grids using hybrid RANS-LES, the highest-fidelity scale-resolving CFD approach routinely deployed by the automotive industry (Ashton et al., 2024). Each sample in the dataset includes a surface mesh with approximately 8.8 million points and corresponding pressure values. Since an official dataset split is not available, we randomly split the dataset into 80% training and 20% testing. We randomly sample 40,000 points in each dataset sample for training and evaluation.
@@
We propose the laser powder bed fusion (LPBF) additive manufacturing dataset which involves field prediction on complex geometries. In metal additive manufacturing, variations in design geometry can affect the dimensional accuracy of the part and lead to shape distortions. We perform several LPBF process simulations of a set of geometries to obtain the deformation field over the geometry. We select a subsample of the dataset with up to 50,000 points per geometry and divide it into 1,100 training and 290 cases. We train models to learn the $Z$ (vertical) component of the deformation field at each grid point. Additional details are provided in Appendix E.
@@ C.3 Benchmark models and training details
In running our experiments, we noticed discrepancies between our LNO results and those presented in their article (Wang & Wang, 2024a). Digging deeper, we found inconsistencies between our training setup that emulates Transolver (code available at https://github.com/thuml/Transolver). and that of LNO (code available at https://github.com/L-I-M-I-T/LatentNeuralOperator) that result in different train/test split. This can be problematic as the authors of LNO make a direct comparison with the results presented in the Transolver paper. We have informed the authors of LNO of this discrepancy and requested further details on their training setup by rasing a pull request at https://github.com/L-I-M-I-T/LatentNeuralOperator/pull/6.
@@
Figure 9: Execution times in FP32 for a single vanilla self-attention layer, a physics attention layer, and FLAME. The models are set to have approximately the same number of parameters as in Section 4.1. This calculation is performed on a single H100 80GB GPU. Note that the curves for FLARE are somewhat overlapping.
@@ Appendix D Field-prediction on million-point geometries
Appendix E Benchmark dataset of additive manufacturing simulations
Figure 10: We simulate the LPBF process on selected geometries from the Autodesk segementation dataset (Lambourne et al., 2021) to generate a benchmark dataset for AM calculations. Several geometries are presented in this gallery. The color indicates $Z$ (vertical) displacement field.
E.1 Introduction & Background
@@
In Ferguson et al. (2025), a dataset of LPBF simulations were released where multiple finite element calculations were carried out on a dataset of 3D shapes. However, that dataset was limited to small, coarse meshes with approximately 3,500 points per mesh. This work is an iteration upon the work of Ferguson et al. (2025) with larger, refined meshes with up to 50,000 grid points.
@@ E.3 Benchmark task
E.4 Data filtering
The dataset contains a wide-varying range of shapes, making the data set general enough to train a strong data-driven field prediction model. Out of $\sim$27,000 shapes, 19,732 were successfully simulated. We analyze the first 3,500 successful simulations and filter them according to several statistics to design a balanced training and test set. For example, we limit the learning problem to meshes with up to 50,000 points and up to 300,000 shapes. This is done to reduce memory usage which becomes a bottleneck when training on small GPUs. We also filter meshes that have high aspect ratio elements as the FEM calculation could be unreliable on highly distorted geometries. Our dataset processing code is available in the GitHub repository associated with this paper. The statistics for the filtered data set are presented in Table 6 along with histograms in Figure 11. A gallery of successful simulations is presented in Figure 10.

### v2

@@ 5 Conclusion
FLARE is a token mixing layer that bypasses the quadratic cost of self-attention by leveraging low-rankness. Mechanically, FLARE routes attention through a fixed-size latent sequence via cross-attention projection and unprojection. FLARE achieves SOTA accuracy on a set of diverse PDE benchmarks, and easily scales to PDE problems with million-scale geometries.
As transformers are the backbone of modern deep learning, we postulate that an efficient attention mechanism has several applications. We also identify potential areas for improvement: FLARE’s reliance on deep residual MLPs can introduce sequential bottlenecks and increase latency, suggesting that further speedups are possible by addressing this issue. Additional enhancements include (1) incrementally increasing the number of latent tokens during training; (2) conditioning latent tokens on time for diffusion modeling; and (3) designing decoder-only variants for auto-regressive modeling.
LLM Usage
LLMs assisted with copyediting (grammar/clarity) and with scaffolding boilerplate in our experimental scripts (e.g., debugging, refactors) inside an IDE with LLM support. LLMs did not generate ideas, architectures, analyses, results, or datasets. All LLM outputs were reviewed, edited, and tested by the authors, who take full responsibility for the content.
@@ C.2 Benchmark datasets
The DrivAerML dataset (Ashton et al., 2024) has high-fidelity automotive aerodynamic simulations, featuring 500 parametrically morphed DrivAer vehicles. CFD simulations are performed on 160 million volumetric mesh grids using hybrid RANS-LES, the highest-fidelity scale-resolving CFD approach routinely deployed by the automotive industry (Ashton et al., 2024). Each sample in the dataset includes a surface mesh with approximately 8.8 million points and corresponding pressure values. Since no official dataset split is provided, an 80/20 random split is used, with 40,000 points sampled per case for training and evaluation.
@@
We introduce the laser powder bed fusion (LPBF) additive manufacturing dataset which involves field prediction on complex geometries. In metal additive manufacturing, variations in design geometry can affect the dimensional accuracy of the part and lead to shape distortions. We perform several LPBF process simulations of a set of geometries to obtain the deformation field over the geometry. We select a subsample of the dataset with up to 50,000 points per geometry and divide it into 1,100 training and 290 cases. We train models to learn the $Z$ (vertical) component of the deformation field at each grid point. Additional details are provided in Appendix G.
@@ C.3 Benchmark models and training details
In running our experiments, we noticed discrepancies between our LNO results and those presented in their article (Wang & Wang, 2024a). Upon further investigation, we found that the datasets used in the LNO paper and in the original Transolver paper are not the same. For example, in the Elasticity dataset, LNO was trained and tested on a $1000/1000$ split, whereas Transolver used a $1000/200$ split. In the Darcy dataset, LNO employed a higher resolution of $241\times 241$, compared to $85\times 85$ in Transolver. Because these differences make direct comparison unreliable, all models (including LNO and Transolver) were re-trained and evaluated on the standardized Transolver splits and resolutions to ensure fairness.
@@
Figure 8: Execution times in FP32 for a single vanilla self-attention layer, a physics attention layer, and FLARE. The models are set to have approximately the same number of parameters as in Section 4.1. This calculation is performed on a single H100 80GB GPU. Note that the curves for FLARE are somewhat overlapping.
@@ Appendix D Field-prediction on million-point geometries
Appendix E Discussion on the design principles of FLARE.
Latent tokens enable gather-scatter communication.
In FLARE, information flows through latent tokens by first gathering from the input sequence and then scattering back. The encoding step can be understood as a gather (all-reduce) operation, where each latent token pools information from the input according to its learned query pattern. Formally, for latent query $q_{m}$,
$z_{m}=\sum_{n=1}^{N}\frac{\exp(q_{m}\cdot k_{n})}{\sum_{n^{\prime}=1}^{N}\exp(q_{m}\cdot k_{n^{\prime}})}\,v_{n},\quad m=1,\dots,M,$ (23)
$z_{m}$ aggregates input values $v_{n}$ with convex weights. When the similarity scores are sharp, $z_{m}$ emphasizes a few dominant inputs; when they are flatter, $z_{m}$ acts like a specialized pooling token that averages a select set of tokens. Across $M$ latents, this yields a compact set of global descriptors, each specializing in pooling different aspects of the input.
The decoding step is the dual scatter (broadcast) operation, but importantly, the broadcast is selective: each latent $z_{m}$ contributes only to the input tokens that assigned it high weight in the encoding step. Concretely,
$y_{n}=\sum_{m=1}^{M}\frac{\exp(k_{n}\cdot q_{m})}{\sum_{m^{\prime}=1}^{M}\exp(k_{n}\cdot q_{m^{\prime}})}\,z_{m},\quad n=1,\dots,N,$ (24)
the output token $y_{n}$ only receives substantial information from the latents whose query pattern matches its key $k_{n}$ strongly. In this sense, each latent acts as both a selective pooling hub and broadcaster, routing information back along the pathways that originally recruited it. Together, these gather–scatter stages form a low-rank, butterfly-like transformation: first contracting $N$ tokens into $M\ll N$ global features, and then expanding back to $N$ outputs, enabling efficient, yet expressive, global communication.
Symmetry in latent token communication.
The compression–expansion attention structure in FLARE is most effective when the encoding and decoding operators are structurally aligned. Specifically, $W_{\text{encode},h}=\mathrm{softmax}(Q_{h}\cdot K_{h}^{T})$ and $W_{\text{decode},h}=\mathrm{softmax}(K_{h}\cdot Q_{h}^{T})$ are transposes of each other up to diagonal scaling. We experimented with breaking this symmetry by using distinct query–key pairs for encoding and decoding, but observed no accuracy gains. This suggests that the near-adjoint relationship between $W_{\text{encode}}$ and $W_{\text{decode}}$, both derived from the same parameters, may provide a form of mathematical optimality, ensuring stable information flow through latent tokens while reducing representational redundancy.
Tradeoff between query dynamics and key/value expressivity.
An important design choice in FLARE is to use fixed, input-independent queries $Q$, which constrain the flexibility of the attention pattern. Detaching $Q$ from $X$ allows for the low-rank communication structure, but requires compensating expressivity in the key and value projections. In practice, we find that deeper residual MLPs for $K$ and $V$ are crucial for capturing rich, feature interactions under this constraint. Conversely, one could imagine making $Q$ dynamic by conditioning it on $X$, which would shift more of the modeling burden to the query side and potentially allow shallower $K,V$ projections. Thus, FLARE embodies a clear tradeoff: fixing $Q$ encourages stability and efficiency, but places greater importance on the depth and expressivity of the key/value encoders.
Appendix F Model analysis and ablations
ResMLP depth: key/value projections.
ResMLP depth: feedforward block.
Figure 10: Effect of head dimension ($D$) on test accuracy. We design FLARE to work optimally for D=4D=488.
Number of heads.
A key hypothesis behind our design is multiple simultaneous (head-wise parallel) low-rank projections collectively approximate a full attention map. Thus, for a fixed total feature dimension ($C$), using a smaller head dimension ($D$) should improve performance by allowing more such parallel projections. To test this, we evaluate FLARE on the elasticity dataset with varying head counts while keeping the total feature dimension $C=64$ and number of blocks $B=8$ fixed. As shown in Figure 10, FLARE achieves the lowest test error when $D=4$ or $D=8$, in contrast to standard transformers which often use D=16D=163232. Notably, because our head dimensions are smaller, we use a scaling factor of $1$ instead of the typical $\sqrt{D}$ employed in transformers (Vaswani et al., 2017), as the dot product magnitudes are sufficiently small without additional scaling.
Appendix G Benchmark dataset of additive manufacturing simulations
Figure 11: We simulate the LPBF process on selected geometries from the Autodesk segmentation dataset (Lambourne et al., 2021) to generate a benchmark dataset for AM calculations. Several geometries are presented in this gallery. The color indicates $Z$ (vertical) displacement field.
G.1 Introduction & Background
@@
Ferguson et al. (2025) introduced a dataset of LPBF simulations in which multiple finite element calculations were performed on a collection of 3D shapes. That dataset, however, was restricted to relatively coarse meshes with approximately 3,500 points per mesh. The present work extends this line of research by considering larger and more refined meshes, with up to 50,000 grid points.
@@ G.3 Benchmark task
G.4 Data filtering
The dataset contains a wide-varying range of shapes, making the data set general enough to train a strong data-driven field prediction model. Out of $\sim$27,000 shapes, 19,732 were successfully simulated. We analyze the first 3,500 successful simulations and filter them according to several statistics to design a balanced training and test set. For example, we limit the learning problem to meshes with up to 50,000 points and up to 300,000 edges. This is done to reduce memory usage which becomes a bottleneck when training on small GPUs. We also filter meshes that have high aspect ratio elements as the FEM calculation could be unreliable on highly distorted geometries. Our dataset processing code is available in the GitHub repository associated with this paper. The statistics for the filtered data set are presented in Table 6 along with histograms in Figure 12. A gallery of successful simulations is presented in Figure 11.